OpenStudy (anonymous):

What volume of carbon dioxide gas at 650 torr and 20°C would be formed from 2 g of sodium bicarbonate? HCl + NaHCO3 → NaCl + CO2 + H2O

OpenStudy (aaronq):

okay, so first things first, the equation should be balanced. Is it?

OpenStudy (anonymous):

give me 2 secs i need to pee sorry I'm so public

OpenStudy (aaronq):

lol kk

OpenStudy (anonymous):

I'm back srry

OpenStudy (aaronq):

no probs. So is the equation balanced?

OpenStudy (anonymous):

i think so? it looks balanced

OpenStudy (aaronq):

it is. i was just checking that you knew that it was as well. so we have to work with the 2g of sodium bicarbonate, since it's all were given (we don't need to find the limiting reactant). So convert the mass to moles, using the molar mass. (divide the mass by the molar mass)

OpenStudy (anonymous):

so its 2g/60g/mol?

OpenStudy (aaronq):

the molar mass of NaHCO3 is 84 g/mol

OpenStudy (anonymous):

so 2g/84g/mol?

OpenStudy (aaronq):

yep. So, like before, we need to set up a ratio with the moles and the coefficients. \(\dfrac{moles~of~HCO_3}{1}=\dfrac{moles~of~CO_2}{1}\) since the coefficients are both the same, then the moles produced are the same.

OpenStudy (anonymous):

so 0.023g/mol/1=moles of co^2/1?

OpenStudy (aaronq):

the units are 0.0238 moles, not g/mol but yeah, that's the correct set up.

OpenStudy (anonymous):

ok so it would just be 0.0238 moles

OpenStudy (aaronq):

so now that you know the moles of \(CO_2\), and since this isn't at STP, use the ideal gas law: PV=nRT

OpenStudy (aaronq):

yeah thats correct.

OpenStudy (anonymous):

so p would be 0.0238 moles?

OpenStudy (anonymous):

i meant v

OpenStudy (aaronq):

nope. P= pressure V=volume (what you're looking for) n=moles R=constant T=temp (in Kelvin)

OpenStudy (anonymous):

and p is 650 torr

OpenStudy (aaronq):

yep, so first identify all your values then just plug them in and solve.

OpenStudy (anonymous):

wats the v?

OpenStudy (aaronq):

V is what you're looking for in the question. Have you determined which is which of the rest of the variables?

OpenStudy (anonymous):

wats r?

OpenStudy (aaronq):

R in this case is 62.364 L torr/ mol·K

OpenStudy (anonymous):

ok

OpenStudy (aaronq):

so plug in all your values and solve for V

OpenStudy (aaronq):

remember that T must be in Kelvin (not celsius)

OpenStudy (anonymous):

do u cancel units before multiplying or after ?

OpenStudy (aaronq):

whenever you want

OpenStudy (anonymous):

so its 11883.29?

OpenStudy (aaronq):

hm thats the volume you got?

OpenStudy (aaronq):

if so, thats not correct. did you divide by the pressure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me do it again?

OpenStudy (anonymous):

28.126?

OpenStudy (aaronq):

\(V=\dfrac{62.364~L ·torr/ mol·K*0.0238~moles*(20+273.15)K}{650 torr}\) is this what you did?

OpenStudy (aaronq):

i got V=0.669402 L = 0.7 L

OpenStudy (anonymous):

i didn't add the 4 on 62.364

OpenStudy (aaronq):

that shouldn't be a big problem though. You should try getting the same answer as me with your calculator

OpenStudy (anonymous):

idk mac calculators are weird

OpenStudy (aaronq):

hm i use this website http://web2.0calc.com

OpenStudy (anonymous):

no i got the same as u my computer was actin weird

OpenStudy (aaronq):

oh ok, good stuff !

OpenStudy (aaronq):

so thats the final answer, the volume of CO2

OpenStudy (anonymous):

do u want me to post the other questions?

OpenStudy (aaronq):

sure, post one at a time though.

OpenStudy (anonymous):

ok